The areaa of a circle is about 28 in2. Te diameter of the circle is about what?
Answer by feliz1965(151)

The area of a circle is about 28in^2.
The diameter of the circle is about what?
=========================================
The area is given to be 28 inches squared.
The words "about what" mean that we are searching for an approximate answer NOT and exact answer.
A = pir^2
28 = pir^2
We need to solve for r. Then we can multiply the answer by 2 because it takes 2 radii to make 1 diameter.
28/3.14 = r^2
2.99 = r
Then diameter = 2r.
diameter = 2(2.99)
diameter = 5.98 inches.
Feliz